# Community-level Contagion among Diverse Financial Assets

### v1

@@ 3.1 Timeseries Description
  • 2.
    Market capitalization: Given thousands of stocks available in the stock market, we restrict our selection to those whose underlying companies rank among the top 200 by market capitalization as listed on Companies Market Cap. For US ETFs and cryptocurrencies, since each category has less than 100 assets available, we consider all of them.
  • 3.
@@ 3.3 Sector Division
To gain a more detailed understanding of the assets and the contagion effects among them, we examine not only their types of investment vehicle but also their business sectors. Specifically, by convention, the stock market is classified into 11 categories based on business sectors, including Communication Services, Utilities, Real Estate, Materials, Information Technology, Industrials, Healthcare, Financials, Energy, Consumer Staples, and Consumer Discretionary. Regarding US ETFs, these are more complex because each represents a basket of multiple stocks and reflects the overall performance of those underlying assets. As a result, a single US ETF may consist of stocks from various sectors. For instance, the well-known S&P500 ETF (e.g. SPY) includes contributions from all sectors, with Technology being the largest one, accounting for 32% of its holdings, as reported by Yahoo!finance. For this characteristic, we assign a US ETF to a specific sector if that sector represents at least 50% of its holdings, otherwise, it is classified as mixed. Cryptocurrencies, on the other hand, are not tied to specific business domains and are therefore treated as a single sector. The distribution of stocks and US ETFs by business sector is shown in Table 2.
Table 2: Classification of stocks (black text) and US ETFs (blue text) by business sector. Only US ETFs with at least 50% of their holdings concentrated in a single sector are classified and considered to belong to that specific sector. Business sectors without available assets are not displayed.
@@
Technology ASX, AAPL, ACN, ADBE, ADI, ADP, AMAT, AMD, ASML, AVGO, BABA, CMCSA, CRM, CSCO, GOOG, IBM, INTC, INTU, LRCX, MA, META, MSFT, MU, NFLX, NOW, NVDA, ORCL, PANW, PDD, QCOM, SAP, SHOP, SONY, TMUS, TSM, TXN, T, VZ, V, CIBR, FNGU, IYW, QQQ, QTEC, QYLD, ROBT, SOXX
@@
Consumer Discretionary ALV, AZMN, BKNG, COST, CSL, DIS, HD, IDEX, LOW, MCD, NKE, NVR, SBUX, TJX, TM, TSLA, ITB
@@ 4.2 Tracy-Widom Theory Recall
However, we cannot adopt the presented theory since our study utilizes the correlation instead of covariance matrix. Therefore, a transformation to adapt to the need of using correlation matrices has been discovered. For this, Edoardo et al. in [75] have introduced another standardization procedure for the eigenvalues obtained from a correlation matrix so that the Tracy-Widom theory is still preserved with the case of correlation matrices. Without loss of generality, we introduce this standardization procedure for the largest eigenvalue. In particular, given the correlation matrix $Cov^{*}$ and its largest eigenvalue $e_{1}^{*}$, the distribution function $D$ of the largest eigenvalue $e_{1}^{*}$ can be experimentally constructed by generating a set of random correlation matrices that have similar characteristics as CovCov* (e.g. mean, standard deviation, size, etc). As a result, there exist parameters $a$ and $b$ such that the standardized eigenvalue $E_{1}^{*}=(e_{1}^{*}-a^{*})/b^{*}$ follows the $F_{1}$ distribution. These parameters were theoretically derived by [87] and have the following forms:
@@
To this end, this theory suggests that the eigenvalues of a random covariance or correlation matrix follow the Tracy-Widom distributions if standardized appropriately. Therefore, it can be used to identify whether an eigenvalue carries informative signals that reflect the underlying characteristics of the correlations between objects or simply carries random noise without real information. In this regard, an eigenvalue is considered to be informative if it falls outside the Tracy-Widom distribution, otherwise, it is considered noise and will be removed [75].
@@ 4.3 Noise Filtering based on Tracy-Widom Theory
Figure 1: Three largest empirical eigenvalues (e.g. $\lambda_{1}>\lambda_{2}>\lambda_{3}$) and their corresponding Tracy-Widom distributions (e.g. TW 1, TW 2, TW 3) for the period from 01/04/2019 to 31/12/2019. All three eigenvalues are informative since they are outside and larger than the Tracy-Widom distributions.
Figure 2: 19th, 20th and 21th empirical eigenvalues (e.g. $\lambda_{19}>\lambda_{20}>\lambda_{21}$) and their corresponding Tracy-Widom distributions (e.g. TW 19, TW 20, TW 21) for the period from 01/04/2019 to 31/12/2019. Although these eigenvalues are outside the Tracy-Widom distribution, they are considered as noise since they are much smaller than the Tracy-Widom distributions, meaning that even if they carry real information, this amount is considered negligible.
From our experiments in this present study, one common result is that the largest eigenvalues carry most of the information content of a correlation matrix, especially $\lambda_{1}$. For instance, this eigenvalue carries more than a third of the total information in the correlation matrix for the Pre-Covid period from 01/04/2019 to 31/12/2019. Notably, there are 221 eigenvalues in total, emphasizing the dominance of largest eigenvalues. As a result, there is not much information left in the following eigenvalues, explaining for the small magnitude of the 19th, 20th and 21th eigenvalues shown in Figure 2.
@@ 4.4 Community Structure Construction
From a correlation matrix, three preprocessing steps need to be conducted to get the community structure of the assets. Specifically, we first transform the correlation values $c_{ij}$ into the distance values $d_{ij}=\sqrt{2\times(1-c_{ij})}$. As a result, the correlation matrix $C$ is transformed into the so-called distance matrix $D_{C}=(d_{ij})_{1\leq i\leq N,1\leq j\leq N}$. This transformation is necessary since the correlation values, although useful in understanding the relationships between entities, do not satisfy the requirements of a metric as defined in Algebra [94]. Therefore, they cannot be used to make quantitative comparisons in terms of distance, which makes them unsuitable for directly evaluating how close or far entities are from each other in a measurable sense. Moreover, such transformation is also mandatory for later algorithms that are only applicable in a metric space.
An asset graph is then constructed from the distance matrix $D_{C}$ where each node of the graph represents an asset and an edge connecting two nodes represents their distance value. A major drawback of the graph lies in its high density where all nodes are connected with each other, making it both memory and computation heavy. Since not all edges contribute equally useful information, we simplify the graph by removing edges with less important information. In particular, our second preprocessing step is using the Kruskal algorithm [95] to extract the Minimum Spanning Tree (MST) [96] of the graph - a subgraph that preserves edges with the highest correlation degree (i.e. the smallest distance values) while maintaining the characteristics of a tree and discarding other edges. This allows us to retain the most important part of the graph while reducing the graph size and redundant information. Consequently, we reduce the graph from 221 nodes and 24,310 edges down to a much sparser structure with 221 nodes and just 220 edges.
@@ 4.5 Contagion Detection
To study how shocks or changes in one time series affect others, the Vector Autoregression (VAR) model is commonly used. Introduced by Sims [97], the VAR model is a statistical tool that helps analyze the relationships between several time series by expressing each variable as a function of its own past values and the past values of the other variables in the group [27]. This allows the model to detect how shocks can move in both directions between any two timeseries. For example, a change in one timeseries may affect another and that second timeseries may in turn influence the first. Given a set of $m$ time series $\{y_{1},y_{2},\cdots,y_{m}\}$, the VAR model of order $p$ can be written as:
@@
where $y_{i}^{t}$ represents the value of time series $y_{i}$ at time $t$. The letter $p$ indicates how many previous time steps (lags) are considered when predicting the current value. Each time series depends not only on its own past values (up to $p$ lags), but also on the past values of all other time series in the system over the same number of lags. The term $c_{i}$ is a constant (intercept) in the equation for time series $y_{i}$. The coefficient $a_{i,j}^{t-z}$ represents the contagion effect of time series $y_{j}$ on $y_{i}$ at lag $z$. If this coefficient is statistically significant at the 90% confidence level or higher, we consider that there is a contagion effect from $y_{j}$ to $y_{i}$ at lag $z$. Furthermore, a positive value means that an increase in one time series tends to be followed by an increase in the other, while a negative value suggests the opposite effect. Finally, $\epsilon_{i}$ is the error term in the equation for $y_{i}$, capturing random fluctuations not explained by the model.
@@ 4.6 Contagion Metrics
We evaluate community level-contagion effects in two ways: overall and directional. The overall effect captures how shocks spread in both incoming and outcoming directions (e.g. bi-directional) among the assets within a community or between two communities. On the other hand, the directional effect measures how contagion flows from one community of assets to another (e.g. unidirectional). Based on these perspectives, we define six metrics outlined below. The design of these metrics is inspired by [98] and [99].
  • 1.
    Overall Contagion Density within a Community($OCD_{individual}$): to measure the density of contagion within a community of assets, considering both incoming and outcoming directions. Let $S$ be a set of contagion links between $n$ assets.
@@
  • 2.
    Overall Contagion Magnitude within a Community ($OCM_{individual}$): to measure the average magnitude of contagion within a community of assets, reflecting how strong the contagion is, on average, within the community. Similar to the $OCD_{individual}$ metric, this considers both incoming and outcoming directions.
@@
  • 3.
    Overall Contagion Density between Two Communities ($OCD_{cross}$): to measure the density of contagion between communities, considering both incoming and outcoming directions. Let $SA$ and $SB$ be the sets of contagion links from $n$ assets in community $A$ to $m$ assets in community $B$ and vice versa, respectively.
@@
  • 4.
    Overall Contagion Magnitude between Two Communities ($OCM_{cross}$): to measure the average magnitude of contagion between two communities, considering both incoming and outcoming directions.
@@ 5 Results and Implications
This section presents and discusses the experimental results on contagion effects during each of the seven aforementioned sub-periods. We recall that this study focuses on contagion at the community level, investigating the phenomenon between communities of assets rather than between individual assets. These communities are identified using the Louvain community detection algorithm, with the Tracy-Widom theory applied to mitigate the impact of noise. Further details of this process are provided in Section 4. The structure of this section follows the two defined research questions and is divided into 4 subsections. First, we provide an overview of the community structure constructed from 221 assets across seven sub-periods. Next, we explore the contagion effects both within individual communities and among each pair of communities. This experiment aims to determine whether intra-community contagion differs significantly from inter-community contagion, answering the first research question. Then, we examine whether certain communities act as dominant transmitters (i.e. transmitting shocks to the rest of the network) or receivers (i.e. receiving shocks from the rest of the network) of contagion signals, thereby addressing the second research question. Finally, implications gathered from the results will be covered in the fourth subsection.
5.1 Community Structures Discovery
We first examine the community structure of 221 assets drawn from 3 categories: stocks, cryptocurrencies and US ETFs, across seven sub-periods, namely Pre-Covid-19, Covid-19 Outbreak, Bull Time 1, Bull Time 2, Bull Time 3, Ukraine-Russia Conflict 1 and Ukraine-Russia Conflict 2. These community structures will be used for experiments in the subsequent subsections. Examples of the community structures are displayed in Figures 3(a) and 4(a) while full details for all sub-periods are provided in the supporting file S2. In these graphs, each node represents an asset and an edge between two nodes indicates a significant similarity (i.e. strong correlation) between them. On the other hand, the absence of an edge suggests a weak or negligible similarity relative to the surrounding connections. Each community is highlighted by a different color. A detailed breakdown of each community within each sub-period is provided in the supplementary file S3.
(a) Community Structure
@@ 5.1 Community Structures Discovery
Figure 3: Community structure (a) and sector distribution per community (b) of 221 assets during the Pre-Covid-19 period. A US ETF is assigned to a specific business sector if that sector accounts for at least 50% of its holdings, otherwise, it is unclassified and labelled as US ETF in the sector distribution.
(a) Community Structure
@@
Figure 4: Community structure (a) and sector distribution per community (b) of 221 assets during the Covid-19 Outbreak period. A US ETF is assigned to a specific business sector if that sector accounts for at least 50% of its holdings, otherwise, it is unclassified and labelled as US ETF in the sector distribution.
@@
Although the community structure changes across sub-periods with assets clustering and dispersing in different ways on a sub-period basis, some patterns remain unchanged overtime. In particular, it can be seen that cryptocurrencies tend to separate themselves from traditional assets, distributing close to each other and forming distinct communities. Exceptions are USDT and DAI - two popular stablecoins designed to track the USD currency, which explains their closer association with traditional markets compared to the cryptocurrency one [100]. By contrast, stocks and US ETFs are more likely to cluster together, emphasizing a strong relationship between them. To support this observation, we measure the average distance between assets within each category and between assets from two different categories, then compare with each other. Specifically, using the Minimum Spanning Tree (MST) and its distance values computed in subsection 4.4, we calculate the pairwise distance between each pair of assets as the sum of the edge weights along the shortest path connecting them. We then compute the average intra-category distances (e.g. stocks versus stocks) and inter-category distances (e.g. stocks versus cryptocurrencies). This results in 6 pairs of comparison, including cryptocurrencies with themselves, stocks with themselves, US ETFs with themselves, cryptocurrencies versus stocks, cryptocurrencies versus US ETFs and stocks versus US ETFs.

### v2

@@ 3.1 Time Series Description
  2. 2.
    Market capitalization: Given thousands of stocks available in the stock market, we restrict our selection to those whose underlying companies rank among the top 200 by market capitalization as listed on Companies Market Cap. For US ETFs and cryptocurrencies, since each category has fewer than 100 assets available, we consider all of them.
  3. 3.
@@ 3.3 Sector Division
To gain a more detailed understanding of the assets and the contagion effects among them, we examine not only their types of investment vehicle but also their business sectors. Specifically, by convention, the stock market is classified into 11 categories based on business sectors, including Communication Services, Utilities, Real Estate, Materials, Information Technology, Industrials, Healthcare, Financials, Energy, Consumer Staples, and Consumer Discretionary. Regarding US ETFs, these are more complex because each represents a basket of multiple stocks and reflects the overall performance of those underlying assets. As a result, a single US ETF may consist of stocks from various sectors. For instance, the well-known S&P500 ETF (e.g. SPY) includes contributions from all sectors, with Technology being the largest one, accounting for 32% of its holdings, as reported by Yahoo!finance. For this characteristic, we assign a US ETF to a specific sector if it is designed to track the performance of stocks exclusively within that sector. Otherwise, it is classified as mixed. Cryptocurrencies, on the other hand, are not tied to specific business domains and are therefore treated as a single sector. The distribution of stocks and US ETFs by business sector is shown in Table 2.
Table 2: Classification of stocks (black text) and US ETFs (blue text) by business sector. Only US ETFs that track the performance of stocks exclusively within a sector are classified and considered to belong to that specific sector. Business sectors without available assets are not displayed.
@@
Technology ASX, AAPL, ACN, ADBE, ADI, ADP, AMAT, AMD, ASML, AVGO, BABA, CMCSA, CRM, CSCO, GOOG, IBM, INTC, INTU, LRCX, MA, META, MSFT, MU, NFLX, NOW, NVDA, ORCL, PANW, PDD, QCOM, SAP, SHOP, SONY, TMUS, TSM, TXN, T, VZ, V, CIBR, IYW, QTEC, SOXX
@@
Consumer Discretionary ALV, AZMN, BKNG, COST, CSL, DIS, HD, IDEX, LOW, MCD, NKE, NVR, SBUX, TJX, TM, TSLA
@@ 4.2 Tracy-Widom Theory Recall
However, we cannot adopt the presented theory since our study utilizes the correlation instead of the covariance matrix. Therefore, a transformation to adapt to the need of using correlation matrices has been discovered. For this, Edoardo et al. in [78] have introduced another standardization procedure for the eigenvalues obtained from a correlation matrix so that the Tracy-Widom theory is still preserved in the case of correlation matrices. Without loss of generality, we introduce this standardization procedure for the largest eigenvalue. In particular, given the correlation matrix $Cov^{*}$ and its largest eigenvalue $e_{1}^{*}$, the distribution function $D$ of the largest eigenvalue $e_{1}^{*}$ can be experimentally constructed by generating a set of random correlation matrices that have similar characteristics as CovCov* (e.g. mean, standard deviation, size, etc). As a result, there exist parameters $a$ and $b$ such that the standardized eigenvalue $E_{1}^{*}=(e_{1}^{*}-a^{*})/b^{*}$ follows the $F_{1}$ distribution. These parameters were theoretically derived by [90] and have the following forms:
@@
To this end, this theory suggests that the eigenvalues of a random covariance or correlation matrix follow the Tracy-Widom distributions if standardized appropriately. Therefore, it can be used to identify whether an eigenvalue carries informative signals that reflect the underlying characteristics of the correlations between objects or simply carries random noise without real information. In this regard, an eigenvalue is considered to be informative if it falls outside the Tracy-Widom distribution. Otherwise, it is considered noise and will be removed [78].
@@ 4.3 Noise Filtering based on Tracy-Widom Theory
Figure 1: Three largest empirical eigenvalues (e.g. $\lambda_{1}>\lambda_{2}>\lambda_{3}$) and their corresponding Tracy-Widom distributions (e.g. TW 1, TW 2, TW 3) for the period from 01/04/2019 to 31/12/2019. All three eigenvalues are informative since they are outside and larger than the Tracy-Widom distributions. In other words, the three empirical eigenvalues lie within the rightmost 1% tail (i.e. rejection area) of their corresponding Tracy–Widom distributions. This rejection area corresponds to the hypothesis that an eigenvalue is informative rather than noise.
Figure 2: 19th, 20th and 21st empirical eigenvalues (e.g. $\lambda_{19}>\lambda_{20}>\lambda_{21}$) and their corresponding Tracy-Widom distributions (e.g. TW 19, TW 20, TW 21) for the period from 01/04/2019 to 31/12/2019. Although these eigenvalues are outside the Tracy-Widom distribution, they are considered as noise since they are much smaller than the Tracy-Widom distributions, meaning that even if they carry real information, this amount is considered negligible. In other words, the three empirical eigenvalues lie outside the rightmost 1% tail (i.e. rejection area) of their corresponding Tracy–Widom distributions.
From our experiments in this present study, one common result is that the largest eigenvalues carry most of the information content of a correlation matrix, especially $\lambda_{1}$. For instance, this eigenvalue carries more than a third of the total information in the correlation matrix for the Pre-Covid period from 01/04/2019 to 31/12/2019. Notably, there are 221 eigenvalues in total, emphasizing the dominance of the largest eigenvalues. As a result, there is not much information left in the following eigenvalues, explaining the small magnitude of the 19th, 20th and 21st eigenvalues shown in Figure 2.
@@ 4.4 Community Structure Construction
From a correlation matrix, three preprocessing steps need to be conducted to get the community structure of the assets. Specifically, we first transform the correlation values $c_{ij}$ into the distance values $d_{ij}=\sqrt{2\times(1-c_{ij})}$. As a result, the correlation matrix $C$ is transformed into the so-called distance matrix $D_{C}=(d_{ij})_{1\leq i\leq N,1\leq j\leq N}$. This transformation is necessary since the correlation values, although useful in understanding the relationships between entities, do not satisfy the requirements of a metric as defined in Algebra [97]. Therefore, they cannot be used to make quantitative comparisons in terms of distance, which makes them unsuitable for directly evaluating how close or far entities are from each other in a measurable sense. Moreover, such a transformation is also mandatory for later algorithms that are only applicable in a metric space.
An asset graph (or network) is then constructed from the distance matrix $D_{C}$ where each node of the graph (or network) represents an asset and an edge connecting two nodes represents their distance value. A major drawback of the graph lies in its high density where all nodes are connected with each other, making it both memory and computation-heavy. Since not all edges contribute equally useful information, we simplify the graph by removing edges with less important information. In particular, our second preprocessing step is using the Kruskal algorithm [98] to extract the Minimum Spanning Tree (MST) [99] of the graph - a subgraph that preserves edges with the highest correlation degree (i.e. the smallest distance values) while maintaining the characteristics of a tree and discarding other edges. This allows us to retain the most important part of the graph while reducing the graph size and redundant information. Consequently, we reduce the graph from 221 nodes and 24,310 edges down to a much sparser structure with 221 nodes and just 220 edges - the ultimate network structure used in this study.
@@ 4.5 Contagion Detection
Although different methods have been developed for contagion detection, the Vector Autoregression (VAR) model remains widely used and has consistently provided trustworthy insights in various studies [52, 53]. Therefore, we choose to use the VAR approach in our present study. Introduced by Sims [100], this is a statistical tool that helps analyze the relationships between several time series by expressing each variable as a function of its own past values and the past values of the other variables in the group [27]. This allows the model to detect how shocks can move in both directions between any two time series. For example, a change in one time series may affect another and that second time series may in turn influence the first. Given a set of $m$ time series $\{y_{1},y_{2},\cdots,y_{m}\}$, the VAR model of order $p$ can be written as:
@@
where $y_{i}^{t}$ represents the value of time series $y_{i}$ at time $t$. The letter $p$ indicates how many previous time steps (lags) are considered when predicting the current value. Each time series depends not only on its own past values (up to $p$ lags), but also on the past values of all other time series in the system over the same number of lags. The term $c_{i}$ is a constant (intercept) in the equation for the time series $y_{i}$. The coefficient $a_{i,j}^{t-z}$ represents the contagion effect of time series $y_{j}$ on $y_{i}$ at lag $z$. If this coefficient is statistically significant at the 90% confidence level or higher, we consider that there is a contagion effect from $y_{j}$ to $y_{i}$ at lag $z$. Furthermore, a positive value means that an increase in one time series tends to be followed by an increase in the other, while a negative value suggests the opposite effect. Finally, $\epsilon_{i}$ is the error term in the equation for $y_{i}$, capturing random fluctuations not explained by the model.
We acknowledge that using a stricter confidence level such as 95% or 99% can provide more rigorous analytical results. However, our goal is to keep as many contagion coefficients as possible to ensure a sufficient sample size within and between communities for statistical testing. Adopting such stricter thresholds would reduce the number of available contagion coefficients, potentially leaving some tests with too few observations and thereby undermining their statistical validity. In contrast, the 90% confidence level remains a widely accepted standard in the literature and allows us to retain a higher number of observations. To this end, we chose to use the 90% confidence level for this analysis.
@@ 4.6 Contagion Metrics
We evaluate community-level contagion effects in two ways: overall and directional. The overall effect captures how shocks spread in both incoming and outgoing directions (e.g. bi-directional) among the assets within a community or between two communities. On the other hand, the directional effect measures how contagion flows from one community of assets to another (e.g. unidirectional). Based on these perspectives, we define six metrics outlined below. The design of these metrics is inspired by [101] and [102].
  1. 1.
    Overall Contagion Density within a Community($OCD_{individual}$): to measure the density of contagion within a community of assets, considering both incoming and outgoing directions. Let $S$ be a set of contagion links between $n$ assets.
@@
  2. 2.
    Overall Contagion Magnitude within a Community ($OCM_{individual}$): to measure the average magnitude of contagion within a community of assets, reflecting how strong the contagion is, on average, within the community. Similar to the $OCD_{individual}$ metric, this considers both incoming and outgoing directions.
@@
  3. 3.
    Overall Contagion Density between Two Communities ($OCD_{cross}$): to measure the density of contagion between communities, considering both incoming and outgoing directions. Let $SA$ and $SB$ be the sets of contagion links from $n$ assets in community $A$ to $m$ assets in community $B$ and vice versa, respectively.
@@
  4. 4.
    Overall Contagion Magnitude between Two Communities ($OCM_{cross}$): to measure the average magnitude of contagion between two communities, considering both incoming and outgoing directions.
@@ 5 Results and Implications
This section presents and discusses the experimental results on contagion effects during each of the seven aforementioned sub-periods, spanning from 01/04/2019 until 03/05/2023. We recall that this study focuses on contagion at the community level, investigating the phenomenon between communities of assets rather than between individual assets. These communities are identified using the Louvain community detection algorithm, with the Tracy-Widom theory applied to mitigate the impact of noise. Further details of this process are provided in Section 4. The structure of this section follows the two defined research questions and is divided into 4 subsections. First, we provide an overview of the network and community structure constructed from 221 assets across seven sub-periods. Next, we explore the contagion effects both within individual communities and among each pair of communities. This experiment aims to determine whether intra-community contagion differs significantly from inter-community contagion, answering the first research question. Then, we examine whether certain communities act as dominant transmitters (i.e. transmitting shocks to the rest of the network) or receivers (i.e. receiving shocks from the rest of the network) of contagion signals, thereby addressing the second research question. Finally, implications gathered from the results will be provided in the fourth subsection.
Although our study covers several major financial and global events, it does not capture the most recent structural changes in traditional and cryptocurrency markets. Given the rapid evolution of market microstructure, regulatory frameworks and financial conditions since 2023, extending the dataset to include more recent observations could provide additional insights and further validate our results. While incorporating such data lies beyond the scope of the present study, we acknowledge this as a limitation and encourage future research to integrate up-to-date information to enhance and extend our findings.
5.1 Network and Community Structures Discovery
We first examine the network and community structure of 221 assets drawn from 3 categories: stocks, cryptocurrencies and US ETFs, across seven sub-periods, namely Pre-Covid-19, Covid-19 Outbreak, Bull Time 1, Bull Time 2, Bull Time 3, Ukraine-Russia Conflict 1 and Ukraine-Russia Conflict 2. These network and community structures will be used for experiments in the subsequent subsections. Examples of the network as well as community structures are displayed in Figures 3(a) and 4(a) while full details for all sub-periods are provided in the supporting file S2. In these graphs, each node represents an asset and an edge between two nodes indicates a significant similarity (i.e. strong correlation) between them. On the other hand, the absence of an edge suggests a weak or negligible similarity relative to the surrounding connections. In a network, each community is highlighted by a different color. A detailed breakdown of each community within each sub-period is provided in the supplementary file S3.
(a) Network and Community Structure
@@ 5.1 Network and Community Structures Discovery
Figure 3: Network and Community structure (a) and sector distribution per community (b) of 221 assets during the Pre-Covid-19 period. A US ETF is assigned to a specific business sector if it is designed to track the performance of stocks within that sector. Otherwise, it is unclassified and labeled as US ETF (i.e. US ETFs that do not track a specific sector or those that include stocks from multiple sectors).
(a) Network and Community Structure
@@
Figure 4: Network and Community structure (a) and sector distribution per community (b) of 221 assets during the Covid-19 Outbreak period. A US ETF is assigned to a specific business sector if it is designed to track the performance of stocks within that sector. Otherwise, it is unclassified and labeled as US ETF (i.e. US ETFs that do not track a specific sector or those that include stocks from multiple sectors)
To quantitatively assess the network structure of assets in each sub-period, we employ three widely used network centrality measures, including Degree Assortativity [103], Betweenness Centrality [104], and Closeness Centrality [104]. In terms of community structures, we also use Betweenness Centrality and Closeness Centrality but on each individual community (i.e. Community Betweenness and Community Closeness) rather than the whole network like the former. We note that the results of each metric will be averaged to get an overall value for each sub-period. While detailed definitions of these measures can be found in the cited references, a brief explanation is as follows: lower values of Degree Assortativity and Betweenness Centrality indicate a more compressed network, where assets are relatively close to one another and tend to form large communities. Conversely, higher values suggest a more distributed structure, with assets further apart and tend to form a larger number of smaller communities. In contrast, the interpretation of Closeness Centrality is reversed: higher values correspond to a more compressed network.
The results of this assessment are presented in Table 3. Overall, the values of each metric on the network structure vary across sub-periods and no single sub-period consistently exhibits the most compressed or most distributed structure. This indicates that the network structure evolves over time and seems not to be influenced by market conditions. At the community level, although the network centrality measures also fluctuate over time, the Covid-19 Outbreak sub-period consistently shows a more compressed structure compared with the others. This indicates that assets became more closely connected within their communities during this time, implying that information flowed more rapidly and efficiently among assets as a result of the market turbulence triggered by the Covid-19 pandemic.
Table 3: Network centrality measures in each sub-period
Network Centrality Measure Pre-Covid-19
Covid-19
Outbreak
Bull Time 1 Bull Time 2 Bull Time 3
U-R
Conflict 1
U-R
Conflict 2
Degree Assortativity -0.435 -0.422 -0.354 -0.368 -0.355 -0.391 -0.397
Betweenness Centrality 0.038 0.043 0.042 0.050 0.048 0.041 0.040
Closeness Centrality 0.127 0.138 0.123 0.101 0.112 0.139 0.135
Community Betweenness 0.139 0.125 0.157 0.160 0.167 0.144 0.150
Community Closeness 0.429 0.516 0.438 0.423 0.457 0.492 0.465
@@
However, although the network and community structure changes across sub-periods with assets clustering and dispersing in different ways on a sub-period basis, some patterns remain unchanged over time. In particular, it can be seen that cryptocurrencies tend to separate themselves from traditional assets, distributing close to each other and forming distinct communities. Exceptions are USDT and DAI - two popular stablecoins designed to track the USD currency, which explains their closer association with traditional markets compared to the cryptocurrency one [105]. By contrast, stocks and US ETFs are more likely to cluster together, emphasizing a strong relationship between them. To support this observation, we measure the average distance between assets within each category and between assets from two different categories, then compare them with each other. Specifically, using the Minimum Spanning Tree (MST) and its distance values computed in subsection 4.4, we calculate the pairwise distance between each pair of assets as the sum of the edge weights along the shortest path connecting them. We then compute the average intra-category distances (e.g. stocks versus stocks) and inter-category distances (e.g. stocks versus cryptocurrencies). This results in 6 pairs of comparison, including cryptocurrencies with themselves, stocks with themselves, US ETFs with themselves, cryptocurrencies versus stocks, cryptocurrencies versus US ETFs and stocks versus US ETFs.